OpenStudy (unofficialllyy):

Can someone check my 10 Geometry questions? I will fan and medal :)

OpenStudy (unofficialllyy):

1. Look at the triangle: What is the value of sin x°? My answer: 5 divided by 13

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

yeah \(\frac{5}{13}\)

OpenStudy (unofficialllyy):

2. Look at the figure: If tan x° = 11 divided by r and cos x° = r divided by s, what is the value of sin x°? My answer: \[\sin x= \frac{ 11 }{ s }\]

jimthompson5910 (jim_thompson5910):

correct on that one too

OpenStudy (unofficialllyy):

3. Max observes the zoo and the library from a helicopter flying at a height of 200 times square root of 3 feet above the ground, as shown below What is the distance between the zoo and the library? My answer: 400 feet

OpenStudy (unofficialllyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (unofficialllyy):

4. If the cos 30° = square root 3 over 2, then the sin 60° = _____. My answer: square root 3 over 2, because the angles are complementary.

jimthompson5910 (jim_thompson5910):

yep sin(x) = cos(90 - x) where 0 <= x <= 90

OpenStudy (unofficialllyy):

5. Use the image below to answer the following question. What relationship do the ratios of sin x° and cos y° share? My answer: The ratios are both identical (6/10 and 6/10)

jimthompson5910 (jim_thompson5910):

correct, same reason as before

jimthompson5910 (jim_thompson5910):

well you can reduce 6/10 to get 3/5

OpenStudy (unofficialllyy):

Two birds sit at the top of two different trees. The distance between the first bird and a bird-watcher on the ground is 14.1 feet. The distance between the bird-watcher and the second bird is 27.9 feet. What is the angle measure, or angle of depression, between this bird and the bird-watcher? My answer: 30.4

OpenStudy (anonymous):

soh cah toa - sin soa - opposite / adajenct enjoy:)

jimthompson5910 (jim_thompson5910):

30.356338108109 degrees which rounds to 30.4 degrees (assuming you are going for the nearest tenth). So you got it

OpenStudy (unofficialllyy):

The picture shows a barn door: What is the length of the support AB? My answer: 9/sin60

OpenStudy (unofficialllyy):

@jim_thompson5910 I only have a few more to be checked :)

jimthompson5910 (jim_thompson5910):

yeah then you use a calculator to compute 9/(sin(60)) make sure you're in degree mode

OpenStudy (unofficialllyy):

The picture shows a triangular island: Which expression shows the value of c? I didn't get this one.

jimthompson5910 (jim_thompson5910):

hint: tan(45) = a/b sin(45) = a/c cos(45) = b/c

OpenStudy (unofficialllyy):

Okay one second :)

OpenStudy (unofficialllyy):

I'm not positive, but I think it's b/cos45

jimthompson5910 (jim_thompson5910):

either that or c = a/sin(45)

OpenStudy (unofficialllyy):

My choices are b/tan45 b/cos45 a cos 45 a tan 45

OpenStudy (unofficialllyy):

So I will go with b/cos45.

OpenStudy (unofficialllyy):

Two ropes, AD and BD, are tied to a peg on the ground at point D. The other ends of the ropes are tied to points A and B on a flagpole, as shown below: Angle ADC measures 60° and angle BDC measures 30°. What is the distance between the points A and B on the flagpole? My answer: 10 feet

OpenStudy (unofficialllyy):

@jim_thompson5910

OpenStudy (unofficialllyy):

I got a 100 :)